OpenStudy (anonymous):

Find the vertex of the quadratic equation: y = 2x^2 + 12x − 4.

OpenStudy (anonymous):

vertex -b/2a -12/4 = -3 y(-3) = 2*(-3)^2 + 12*(-3) - 4 = 18 - 36 - 4 = -22 (-3,-22)